# Skeleton-OOD: An End-to-End Skeleton-Based Model for Robust Out-of-Distribution Human Action Detection

### v2

@@ 4.3.7 Detection result analysis and visualization
OOD classes hand waving, kicking something, rub two hands, neck pain, touch pocket
It displays the count of samples and their class names respectively from OOD samples that are misclassified into ID classes. Upon examining the class names, it is apparently that certain actions have subtle distinctions but still similar, such as ‘clapping’ and ‘hand waving’, or ‘pat on back’ and ‘neck pain’ as shown in Figure 7. How the model captures the little differences between these actions still needs some improvement.
@@ 5 Conclusion and future works
Admittedly, the model still has some potential limitations. Firstly, since the model is trained on GCN, its generalization ability is constrained by the number of skeleton nodes across various datasets. In real-world scenarios, if the trained model is deployed to identify whether a test sample with a different number of nodes is out-of-distribution, its accuracy may be compromised. Secondly, the end-to-end model’s generalization capability is still bounded by the model parameters and the size of the dataset. The way the activation method is used during training causes some of the improvement in results to come from the training operation. One potential solution is to explore semi-supervised learning techniques to mitigate this issue. Moreover, the problem can be further addressed by introducing generative model and conducting training on datasets with limited ID samples. By leveraging existing category text and feature information, the model can infer the name of the unknown category for unfamiliar sample features. In summary, human action recognition tasks present diverse application scenarios, with numerous novel challenges awaiting exploration based on specific practical contexts.

### v3

@@ 4.3.7 Detection result analysis and visualization
OOD classes hand waving, kicking something, rubbing two hands, neck pain, touch pocket
Table 10 lists out the count of samples and their class names respectively from OOD samples that are misclassified into ID classes. Upon examining the class names, it is apparent that certain actions have subtle distinctions but are still similar, such as ‘clapping’ and ‘hand waving’, or ‘pat on back’ and ‘neck pain’ as shown in Figure 7. How the model captures the little differences between these actions still needs some improvement.
@@
4.3.8 Randomly joint masking result
In order to test the scalability of the model on multiple tasks, we also considered the results of the recognition and classification problems with missing node information in reality. The results of ID and OOD detection on the NTU60 data are shown in Table 11. And plot the trend of AUROC and TOP1 results as shown in Figure 8.
Table 11: Test results of different joint mask percentages (p). Using our model (ASH-P), we randomly masked different percentages of joint information on the NTU-RGB+D 60 ID-OOD mix dataset.
OOD ID
ERROR ↓ FPR95 ↓ AUROC ↑ Top1 ↑ Overlap ↓
origin(0) 30.42 52.68 86.64 91.94 0.39
p=5% 33.99 52.32 73.97 47.29 0.46
p=7% 33.56 53.27 73.62 46.58 0.46
p=10% 33.74 52.86 75.30 46.06 0.49
p=15% 33.72 52.91 75.51 45.19 0.49
p=20% 33.98 52.33 76.06 44.82 0.49
p=30% 34.18 51.90 76.03 44.05 0.49
p=50% 34.11 52.04 75.34 43.91 0.49
p=70% 34.32 51.60 75.36 43.88 0.49
From the results we can clearly see that randomly masking on joints affects the accuracy of the ID data most, but slightly on OOD samples. Especially when the percentage is small, the OOD result is not just a simple downward trend. This may be related to the position of the nodes randomly masked.
Figure 8: Results of ID and OOD performance under different masking percentages.
@@ 5 Conclusion and future works
Admittedly, the model still has some potential limitations. Firstly, since the model is trained on GCN, its generalization ability is constrained by the number of skeleton nodes across various datasets. In real-world scenarios, if the trained model is deployed to identify whether a test sample with a different number of nodes is out-of-distribution, its accuracy may be compromised. Secondly, the end-to-end model’s generalization capability is still bounded by the model parameters and the size of the dataset. The way the activation method is used during training causes some of the improvement in results to come from the training operation. One potential solution is to explore semi-supervised learning techniques to mitigate this issue. Moreover, the problem can be further addressed by introducing generative models and conducting training on datasets with limited ID samples. By leveraging existing category text and feature information, the model can infer the name of the unknown category for unfamiliar sample features. In summary, human action recognition tasks present diverse application scenarios, with numerous novel challenges awaiting exploration based on specific practical contexts.